FLOSS01: Logbook Entries

FLOSS01: Radiometers Messages: 15 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
178 Wed 27-Mar-2002Radiometers and soil sensors removed from all sitesall stnsdelany
172 Wed 13-Mar-2002rad cleanedtoweroncley
164 Mon 04-Mar-20024-component aspiration fans disconnected 3/2stn3militzer
148 Thu 21-Feb-2002Radiometer visitsall stnsdelany
138 Mon 18-Feb-2002rads cleanedall stnsoncley
106 Fri 04-Jan-2002Radiometer Domes cleanedall stnsmilitzer
99 Wed 26-Dec-2001Rsw.in(main) oddtoweroncley
91 Mon 17-Dec-20014-Component Radiometer Stn2 moved to Stn3stn2militzer
89 Thu 13-Dec-2001calibration files for CR10Xsall stnssemmer
86 Mon 10-Dec-2001Change of sw.out at maintowerdelany
83 Mon 10-Dec-2001level sensors temperature errortoweroncley
81 Sun 09-Dec-2001rads cleaned?stn2oncley
78 Sun 09-Dec-2001rads servicedtoweroncley
59 Thu 06-Dec-2001radiometers and soil sensorsall stnsdelany1
56 Thu 06-Dec-2001checked Rsw.outtoweroncley


178: Radiometers, Site all stns, Wed 27-Mar-2002 09:41:17 MST, Radiometers and soil sensors removed from all sites
A special trip was made to FLOSS to recover the radiometers and soil probes
Tony, Charlie and Pierce removed the sensors at Bare, Sage and then Tower

(Also the EVE boxes were removed at the bare and sage sites - so no
data from those sites after 15:00 MST March 26.   GDM)

Bare 12:15
Ground not frozen and soil sensors easily uprooted
Net Q97035
Fluxplate H993563
Tsoil NCAR 0001
Soil moisture 013
Soil sensors placed in Netrad box

Campbell 13527

Sage 12:30 
The up looking Pyrg and K&Z were each shadowed using Steve O's wand, for
three successive periods of 30 s at ~ 12:40
Net Q94125
Flux
SoilT
SoilQ

Campbell 11173

Tower 13:30
The up looking Pyrg and K&Z were each shadowed using Steve O's wand, for
three succesive 30 s periods, for three 10 s periods and for one 60s period


Net Q99258
Flux
SoilT
SoilQ

Campbell 11174

The CO2 cal gases: 
#6 Local 0/0
#5 Hical 700/20


172: Radiometers, Site tower, Wed 13-Mar-2002 16:10:37 MST, rad cleaned
Shortly after the krypton cleaning about 30min ago, we also cleaned the pygs and psps (but I forgot net!).

164: Radiometers, Site stn3, Mon 04-Mar-2002 11:29:04 MST, 4-component aspiration fans disconnected 3/2
4-Component FANS Disconnected:  3/2/02 ~16:30z
to reduce the power consumption.

The dome ventillator fans were pulling too much current
when the temperature was dropping too low.
There will probably be no measurable effect in doing
this because of minimal dust and ample winds in
North Park during this time of the year.

Note this was also entered in the 'station outage'
log entry.
148: Radiometers, Site all stns, Thu 21-Feb-2002 12:38:40 MST, Radiometer visits
At 09:45 visited Sage. No frost on radiometers but some on frost detector and 
on top dome of net radiometer. Cleaned frost from frost detector and net 
radiometer. Handkerchief placed over frost detector for 10 minutes.
All fans operating. Power dissapation resister cold.

At 10:08 visited tower site. No frost on radiometers nor frost detector. Net
radiometer had frost on upper dome. Covered uplooking frost detector from 10:09
until 10:30. Covered downlooking frost detector from 10:30 til 10:40.

138: Radiometers, Site all stns, Mon 18-Feb-2002 15:46:12 MST, rads cleaned
All rads have now been cleaned.
13:55 local - main tower
~15:00 local sage site
~15:15 local bare site

106: Radiometers, Site all stns, Fri 04-Jan-2002 15:57:35 MST, Radiometer Domes cleaned
12:45 Radiometers on Tower, all 6 fans OK, Rnet level OK.

15:30 Radiometers on Sage-brush site, all fans OK, Rnet level OK,

On both sites the domes appeared clean before wiping off.

On both sites there was some snow on top of the up-looking radiometers shield.
99: Radiometers, Site tower, Wed 26-Dec-2001 12:02:22 MST, Rsw.in(main) odd
Rsw.in at the main tower has a lot of extra variation at night compared to sage
and compared to both Rsw.out.  This has been going on for a long time -- perhaps
since the beginning.  Could its fan be bad?

91: Radiometers, Site stn2, Mon 17-Dec-2001 17:26:43 MST, 4-Component Radiometer Stn2 moved to Stn3
Q7 installed in place of 4-component.
Soil Sensors remain.
Coefficients for Q7 changed at about 00:20Z 18Dec01
Q7 at 1.97m oriented south on guyed T-post.
89: Radiometers, Site all stns, Thu 13-Dec-2001 10:45:06 MST, calibration files for CR10Xs
Below yare the calibration coefficients used in
each CR10X.
NOTE: Before the installation of stn3 at the "sage" site, stn2 at "bare"
had the 4 comp system. This system was moved to stn3.
The soil sensors and the Q7 stayed at stn2 and a new
set of soil sensors and Q7 along with the 4 comp were
installed at stn3. So if you are looking at data for
stn2 before the installation of stn3, you need to use
the 4 comp coefficients listed in stn3.

TOWER

sensor	SN	     location	calibration
Q7	Q99258	A+	00	9.35
		A-	01	11.55
PYGIN	26416		02	280.1120448
PYGOUT	29260		03	311.5264798
KZIN	940187		04	100.7049345
KZOUT	940186		05	82.98755187
HFT	H983212		06	38.4
PIcase			10	2500
		A0	11	0.00335513
		A1	12	0.278704
		A2	13	1.4317
		A3	14	998.718
PIdome			15	2500
		A0	16	0.00305339
		A1	17	0.268816
		A2	18	8.77845
		A3	19	-2267.84
POcase			20	2500
		A0	21	0.00335427
		A1	22	0.279319
		A2	23	5.19638
		A3	24	-247.625
POdome			25	2500
		A1	26	0.00305418
		A2	27	0.268098
		A3	28	4.06691
		A4	29	403.649
Tsoil	NCAR002	A0	30	-241.477805
		A1	31	231.052081
		A2	32	9.262642
level  N0383 X, A0	33	0.05750311
		A1	34	14.26792000
		A2	35	-0.13395810
		A3	36	-0.20585990
		A4	37	0.0000000
		A5	38	0.0000000
	     Y, A0	39	-0.02092993
		A1	40	14.44724000
		A2	41	0.12732100
		A3	42	-0.39985070
		A4	43	0.0000000
		A5	44	0.0000000
CS615	12	A0	45	-0.187
		A1	46	0.037
		A2	47	0.335
ice on/off 1=on, 0=off	50	1
min bat	(volts)		51	11
X time	(secs)		52	300
outmode	0=a, 1=b	55	0


stn2 at bare site
sensor	SN	     location	calibration	
Q7	Q97035	A+	00	9.26	
		A-	01	11.52	
HFT	H993563		06	38.9	
Tsoil	NCAR001	A0	30	-241.6249	
		A1	31	230.9534	
		A2	32	9.0362	
CS615	13	A0	45	-0.187	
		A1	46	0.037	
		A2	47	0.335	

stn3 at sage site
sensor	SN	     location	calibration   (sensors at stn2 before stn3)
Q7	Q94125	A+	00	8.64			
		A-	01	10.78
PYGIN	31978		02	280.8594299		*
PYGOUT	31976		03	259.6053998		*
KZIN	940185		04	80.25682183		*
KZOUT	970380		05	86.50519031		*
HFT	H993560		06	37.7
PIcase			10	2500			*
		A0	11	0.00335405		*
		A1	12	0.28145			*
		A2	13	3.04961			*
		A3	14	289.824			*
PIdome			15	2500			*
		A0	16	0.0030332		*
		A1	17	0.291681		*
		A2	18	-3.4779			*
		A3	19	870.025			*
POcase			20	2500			*
		A0	21	0.00335116		*
		A1	22	0.285883		*
		A2	23	0.220836		*
		A3	24	756.272			*
POdome			25	2500			*
		A1	26	0.00302872		*
		A2	27	0.297063		*
		A3	28	-5.33875		*
		A4	29	1070.53			*
Tsoil	NCAR003	A0	30	-242.721555
		A1	31	233.953195
		A2	32	8.282468
level  N2018 X, A0	33	-0.3555644		*
		A1	34	11.0103000		*
		A2	35	-0.1015489		*
		A3	36	-0.4473981		*
		A4	37	0.0000000		*
		A5	38	0.0000000		*
	     Y, A0	39	0.5774926		*
		A1	40	10.6420400		*
		A2	41	0.0040844		*
		A3	42	-0.1290564		*
		A4	43	0.0000000		*
		A5	44	0.0000000		*
CS615	8	A0	45	-0.187
		A1	46	0.037
		A2	47	0.335
ice on/off 1=on, 0=off	50	0
min bat	(volts)		51	0
X time	(secs)		52	0
outmode	0=a, 0=b	55	1



86: Radiometers, Site tower, Mon 10-Dec-2001 13:36:09 MST, Change of sw.out at main
Between 11:50 - 12:10 at the main radiation site the down-looking K&Z was 
swapped out
K&Z 940186 and its internal to the housing cable was removed.
K&Z 940181 and its associated cable was installed.

All domes were cleaned, all fans were checked and the Q7 checked for level.
The frost sensors were covered for a few minutes

Ninety minutes later the radiometer data were viewed. The new radiometer shows adata stream comparable to the corresponding data stream at the Bare Eve site.
83: Radiometers, Site tower, Mon 10-Dec-2001 09:31:30 MST, level sensors temperature error
It is obvious that the level sensors at the main site respond inversely to
temperature -- about 1 angle degree to -0.03 C.  Both X & Y show this behavior.

The little bit of data that I have from the bare site don't show this effect.
81: Radiometers, Site stn2, Sun 09-Dec-2001 17:03:53 MST, rads cleaned?
I attempted to clean the bare site radiometers about 1640 using cotton swabs.
Both pygs had a fair amount of dirt, especially near the base of the dome.
I also cleaned some obvious, though small, streaks on the top psp.

Note: Radiometers should be cleaned in daylight (preferably nearly clear),
so that the effect of cleaning can be quantified.  I realized too late that
it would have been better to do this cleaning in more sun.

Also note that it is hard to do clean radiometers (especially psp's) well with
a cotton swab.  We should have some clean wipes here.
 
78: Radiometers, Site tower, Sun 09-Dec-2001 11:14:41 MST, rads serviced
I opened up Rsw.out to see if I could see a bad connection, but everything 
looked okay.  It was opened about 9:30 and closed about 11:00 and the ladder
was underneath, so this morning's data are bad.

I also attempted to clean Rsw.out with distilled water, but it immediately 
froze.  There still was ice on it when I left.  The others I simply wiped with
a cotton swab.
59: Radiometers, Site all stns, Thu 06-Dec-2001 12:25:28 MST, radiometers and soil sensors
Radiometers and soil sensors

Radiation stands were set up at the Main site and the Bare site on 28 Dec''01

The radiometers were installed at Bare on 29 Nov''01
The radiometers and snow sensor were installed at Main on 30 Nov''01

The soil sensors were installed at Main on 30 Nov''01
The soil sensors were installed at Bare on 30 Nov''01

The radiometers were leveled and fan function were checked at Main on 01 Dec''01
The radiometers were leveled and fan function were checked at Bareon 01 Dec''01

The complement of sensors is shown in ...../doc/radioms.idr (to be copied from
~/deploymnets/floss02/radioms.idr)

56: Radiometers, Site tower, Thu 06-Dec-2001 09:39:42 MST, checked Rsw.out
We noticed (belatedly) last night that Rsw.out.main has been bad.  I reseated
cables this morning about 9:20 and did two tests about 9:30- 20 seconds with
my hand underneath followed by 20 seconds with the guy tension meter wood box
underneath.  I'll look at these data soon.

With my hand, the value was 22 W/m2, with the box 18 W/m2, uncovered about
70 W/m2.  Thus, it appears to be working now.  I hope it stays that way.